OpenStudy (anonymous):

Rewrite sqrt(1+cos 45/2 using a half-angle identity. sin 22.5° cos 22.5° sin 90° cos 90°

OpenStudy (anonymous):

since this time you have to use those formulas, we can't use arctan this time :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tan^(-1)(x)?

OpenStudy (anonymous):

I hate this sht

OpenStudy (anonymous):

no, you have to take the right side of the half angle formula and basically come up with the angle

OpenStudy (anonymous):

\[\cos \frac{ a }{ 2 } = \sqrt{\frac{ 1+\cos(a) }{ 2}}\]

OpenStudy (anonymous):

where a is the angle

OpenStudy (anonymous):

\[\sin \frac{ a }{ 2 } = \sqrt{\frac{ 1-\cos(a) }{ 2 }}\]

OpenStudy (anonymous):

It gives you the angle, so which one is it given the formula?

OpenStudy (anonymous):

trying to plug them in my calculator

OpenStudy (anonymous):

ins't it sin 22.5?

OpenStudy (anonymous):

If it was sin, would there be a + or - in the formula?

OpenStudy (anonymous):

so its cos 22.5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

this is a better website, http://www.trans4mind.com/personal_development/mathematics/trigonometry/HalfAngles.htm

OpenStudy (anonymous):

The first one I posted had too many things going on for even me to want to look at

OpenStudy (anonymous):

Which expression is equivalent to cos(4x) + cos(2x)?

OpenStudy (anonymous):

2cos(3x)cos x -2sin(3x)sin x 2sin(3x)sin x -2cos(3x)cos x

OpenStudy (anonymous):

Do you know the double angle formula?

OpenStudy (anonymous):

yea it does have to many formulas gives me head ache just looking at it

OpenStudy (anonymous):

4x is just a double double angle :P

OpenStudy (anonymous):

you want me too open a new question and post it their so you will get your medals?

OpenStudy (anonymous):

I don't really care. Do it because I actually have to do something really fast

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so someone else will answer when I'm gone